find (a) magnitude of the electric field 0.5m from a 2.5 -μC point charge and (b) the magnitude and direction of the electrostatic force acting on an electron placed at that point.

1
Expert's answer
2022-02-15T10:40:50-0500

Given:

"r=\rm 0.5\: m"

"q=2.5\:\rm μC=2.5*10^{-6}\: C"


The magnitude of electric field

"E=k\frac{q}{r^2}=\rm 9*10^9*\frac{2.5*10^{-6}\: C}{(0.5\: m)^2}=90000\: V/m"

The electrostatic force acting on an electron placed at that point

"F=eE\\
={\rm 1.6*10^{-19}\: C*90000\: V/m}\\
=1.44*10^{-14}\:\rm N"

The force is directed toward point charge.